PLZ help me with this homework and plz HANDWRITE IN A TEXT! not in PIC since its hard to read from. Thank you! (electronic graph!)!!

1. Demand is highest for watermelon in the summer, yet that is also when prices are lowest. Draw a graph showing both demand and supply for watermelon in both the summer and the winter (i.e. two demand curves and two supply curves on one graph) that illustrates how this situation could be possible.

  1. The market for high-quality, one-hour impressions of Cher can be described using the demand and supply equations listed below. Graph this market, and determine 1) the market clearing price (P), 2) the market clearing quantity, and 3) total economic surplus when the market clearing price is charged.

QD = 5000 – 100P

QS = -1000 + 50P
Answer #1

(1)

In equilibrium, QD = QS

5000 - 100P = - 1000 + 50P

150P = 6000

P = 40

(2)

Q = 5000 - 100 x 40 = 5000 - 4000 = 1000

(3)

From demand function: When QD = 0, P = 5000/100 = 50 (vertical intercept) & when P = 0, QD = 5000 (Horizontal intercept)

From supply function: When QS = 0, P = 1000/50 = 20 (vertical intercept)

Therefore,

Economic surplus = (1/2) x (50 - 20) x 1000 = 500 x 30 = 15000

GRAPH:

Data table used:

P QD QS
0 5000 -1000
10 4000 -500
20 3000 0
30 2000 500
40 1000 1000
50 0 1500

Graph: